OpenStudy (anonymous):

Given the following cards, what is the probability of drawing a blue card, not replacing it, and then drawing another blue card?

OpenStudy (anonymous):

3/10 3/20 3/25

OpenStudy (anonymous):

I have no clue sorry.. I dont get the replacing thing

OpenStudy (anonymous):

ok

OpenStudy (zarkon):

\[\frac{3}{5}\times\frac{2}{4}\]